OpenStudy (anonymous):

using elimination cancellation solve the problem for x and y 5x-6y=48 2x=5y=-3

OpenStudy (anonymous):

\[5x-6y=48[1]\] \[2x+5y=-3[2]\] \[[1] \times 2\] \[2(5x-6y)=2(48)\] \[10x-12y=96[3]\] \[[2] \times 5\] \[5(2x+5y)=5(-3)\] \[10x+25y=-15[4]\] \[[3]-[4]\] \[(10x-12y)-(10x+25y)=(96)-(-15)\] \[-12y-25y=96--15\] \[-37y=96+15\] \[-37y=111\] \[y=-3\]

OpenStudy (anonymous):

Find x by replacing y with -3 in one of the equations,

OpenStudy (anonymous):

thanks